OpenStudy (anonymous):

Patty's Pet Company charges $0.55 per pound to ship pet supplies. Part A: Write an equation to determine the total cost, c, to ship p pounds of pet supplies. Use your equation to determine the cost of shipping 2 pounds of pet supplies. Part B: If the company reduces the cost to ship pet supplies by 0.05 per pound, write an equation to determine the total cost, c, to ship p pounds of pet supplies with the reduced cost.

OpenStudy (anonymous):

@Michele_Laino @confluxepic @sleepyjess @brucebaner

OpenStudy (anonymous):

Does anyone know how to do this???

OpenStudy (michele_laino):

I'm trying.....

OpenStudy (michele_laino):

I think it is: c=0.55*p so for two pounds, we have: c=0.55*2=1.10$

OpenStudy (anonymous):

thx sooooo much for explaining part a but what about part b??

OpenStudy (michele_laino):

second part: I thinkl it is: c=(0.55-0.05)*p so c=0.50*p

OpenStudy (anonymous):

thanks @Michele_Laino! I gave you a medal :D

OpenStudy (michele_laino):

thanks!!! @77777jeannie77777

OpenStudy (anonymous):

oh and could you help me with some more?? I'll open another question

OpenStudy (anonymous):

ur welcome :D

OpenStudy (michele_laino):

yes sure!

OpenStudy (michele_laino):

thanks!!!!